Find the complete solution of:



1.) (x ^ 2 y + y ^ 3 - y) d x + (x ^ 3 + xy ^ 2 + x) dy = 0

1
Expert's answer
2022-03-07T10:46:02-0500
"(x^2y+y^3-y)dx+(x^3+xy^2+x)dy=0\\
(x^2y+y^3-y)dx=-(x^3+xy^2+x)dy\\
(1/3+1)d[x^3y+x(y^3-y)]=-(1/3+1)d[xy^3+y(x^3+x)]\\
d[x^3y+x(y^3-y)]=-d[xy^3+y(x^3+x)]\\

\intop{d[x^3y+x(y^3-y)]}=\intop{-d[xy^3+y(x^3+x)]}\\
x^3y+x(y^3-y)=-xy^3+y(x^3+x)+C, C-const\\
x^3y+xy^3-xy=-xy^3+x^3y+xy+C\\
2xy^3-2xy=C\\
xy^3-xy=C/2=C_1, C_1-const\\
x(y^3-y)=C_1\\
x=C_1/(y^3-y)"

Answer: "x=C_1\/(y^3-y), \u0421_1-const" and (x=0,y=0) is also a solution.